# LayerShuffle: Enhancing Robustness in Vision Transformers by Randomizing Layer Execution Order

### v1

@@ Abstract
Due to their architecture and how they are trained, artificial neural networks are typically not robust toward pruning, replacing, or shuffling layers at test time. However, such properties would be desirable for different applications, such as distributed neural network architectures where the order of execution cannot be guaranteed or parts of the network can fail during inference. In this work, we address these issues through a number of proposed training approaches for vision transformers whose most important component is randomizing the execution order of attention modules at training time. We show that with our proposed approaches, vision transformers are indeed capable to adapt to arbitrary layer execution orders at test time assuming one tolerates a reduction (about 20%) in accuracy at the same model size. We also find that our trained models can be randomly merged with each other resulting in functional (”Frankenstein”) models without loss of performance compared to the source models. Finally, we layer-prune our models at test time and find that their performance declines gracefully.
@@ 1 Introduction
While demonstrating impressive performance in many domains (e.g. see Krizhevsky et al. (2012); Vaswani et al. (2017); Radford et al. (2021); Rombach et al. (2022)), deep learning systems demand both extensive computational resources and tight integration of their parts. For applications at scale, they therefore increasingly require the construction of large datacenters with thousands of dedicated hardware accelerators. A paradigm shift from central to decentral model inference, where loosely coupled neural networks are distributed over a number of edge devices that share the computational load (Gacoin et al., 2019) of the model therefore seems ultimately desirable. Unfortunately, current deep learning models lack the robustness necessary for such a paradigm shift.
In general, artificial neural networks (ANNs) (Schmidhuber, 2022) are not robust toward pruning or replacing network layers during deployment. Similarly, changing the order of execution in-between layers without further training usually results in catastrophic losses in accuracy. Nevertheless, these properties would be desirable e.g. in distributed setups as described above, where a model is executed on a number of shared nodes in a network. In such a setting, overloaded or malfunctioning nodes could simply be skipped in favor of other available nodes. Furthermore, malfunctioning nodes or absent nodes could simply be replaced by a similar (not the same) node, allowing for simple logistics when deploying models in practice.
@@
The more recently discovered transformer (Vaswani et al., 2017) architecture has been shown to be more flexible. Transformers, when trained accordingly, can be layer-pruned at test-time (Fan et al., 2019), and recent work merges similar transformer-based language models (Akiba et al., 2024), all with only moderate reduction or even an improvement in performance. We hypothesize that the reason for the high adaptability of transformers can be found in self-attention modules being able to adapt their output based on the received input and that it should be possible to train a transformer network to not only adapt to the variation of its input features based on the overall network input but also the variations caused by receiving input from different layers during test time.
We propose and evaluate three training approaches for vision transformers to address the robustness issues laid out above. The most important component common to all approaches is randomizing the execution order of the vision transformer’s stacked self-attention-and-feed-forward modules at training time (Figure 1(a)). More precisely, we have the following main contribution in this paper:
  • We find that using our training approaches, the layers of a vision transformer are indeed capable of adapting to an arbitrary execution order at test time, assuming one tolerates a moderate reduction in performance. Providing each layer additionally with its current position in the network improves performance only slightly compared to a model without it, suggesting that each attention layer is already capable of determining its role based on the incoming data alone.
  • We perform a UMAP analysis that reveals that layers of models trained with our approach adjust their output depending on which position they hold in the network.
  • We find that trained models can be layer-pruned at test time similar to the models trained with the techniques proposed in Fan et al. (2019), where their performance declines gracefully, i.e. models with reduced amounts of layers still remain functional.
  • In addition, vision transformers, which have been made robust to execution order, can be merged with each other resulting in merged (”Frankenstein”) models without loss of performance compared to the source models.
@@ 2 Related work
Zhu et al. (2020) find that for particular subsets of inputs, transformers perform better when changing the execution order of layers to an input-specific sequence. They optimize the execution order per sample in order to maximize the performance of the model for natural language processing tasks. While the goal in their work is to find a layer sequence of a pre-trained model that is optimal for a given input, our approach aims to make the model robust to any sequence of execution, where layers might even be missing.
In parallel to our work on vision transformers and within the wider framework of mechanistic interpretability for large language models (LLMs), Lad et al. (2024) found that LLMs are very robust to changing the positions of adjacent layers or ablating single layers from the model. The main difference to our work is that the authors do not perform any refinement on the models and switch and ablate layers locally with the aim of better understanding the inner workings of LLMs. On the other hand, we focus on methods and training approaches to increase this innate robustness of the transformer architecture to a point where models at test time function regardless of their layer execution order, and respond gracefully to the ablation of several layers in any stage of the network.
@@
Recent work improves the performance of LLMs on predefined tasks, by merging them using evolutionary strategies (Akiba et al., 2024). Similar to Zhu et al. (2020), the authors’ overall aim is to increase performance rather than robustness in distributed environments, so contrary to our approach, layer execution order and scaling for reduced numbers of layers are in general not considered.
Work on introducing permutation invariance into neural networks has been conducted by Lee et al. (2019), Tang & Ha (2021) as well as Pedersen & Risi (2022). The corresponding former two approaches exploit the permutation equivariance of attention, i.e. the fact that the order in which a sequence of vectors gets presented to the attention module does not change its result, but merely shuffles the sequence of output vectors. This equivariance is achieved by using a fixed-seed query vector in order to obtain an permutation invariant latent code. This latent code stays the same no matter in which order input tokens/patches are presented to the module. The main contrast to our work here is that we exploit permutation invariance in the order of layer executions rather than the order of tokens and patch embeddings and can therefore not make use of permutation equivariance of the attention operation, as it does not apply to switching inputs and outputs. Pedersen & Risi (2022) also consider permutation invariance with respect to the input, although the emphasis of their approach lies on a number of smaller recurrent neural networks Werbos (1988); Hochreiter et al. (2001) with shared weights, which extract a permutation-invariant representation of the input vector in a self-organized way without utilizing the permutation-equivariance of self-attention modules.
@@ 3 Methods
After giving a brief overview on vision transformers (Section 3.1), we investigate three approaches to them for arbitrary layer execution order: First, we simply permute the order of layers randomly during training, such that every training batch is presented to the network’s layers in a different random order (Section 3.2). Second, while randomly permuting the layer order as in the previous approach, we use an layer-depth encoding inspired by learned word embedding approaches (Section 3.3) to see if this additional information would further improve performance. Third, while randomly permuting layer order as in the previous approaches, we try to predict from the output of every layer at which position the layer is currently located in the network using a small layer position prediction network for every layer (Section 3.4). We now describe these approaches in more detail.
3.1 Vision Transformers
Dosovitskiy et al. (2020) have successfully adapted the transformer architecture to computer vision by introducing a preprocessing step that converts images to suitable sequences. They do so by splitting an image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ into a sequence of N flattened patches $\mathbf{x_{p}}\in\mathbb{R}^{N\times(P^{2}C)}$, and then pass each patch through a linear embedding layer $\mathbf{E}\in\mathbb{R}^{(P^{2}C)\times D}.$ $H$,$W$ and $C$ are here the height, width and number of channels of the image respectively and $P$ is the patch size. $D$ is the internal latent dimension of the transformer which remains constant throughout the network and can be set as a hyperparameter.
After converting the image into a sequence that can be processed by a transformer encoder, inspired by BERT (Devlin et al., 2018), the authors prepend a class token to $\mathbf{x_{p}}$ in which the class information of the input image can be aggregated by the transformer. To encode position information into the embedding, a positional embedding tensor $\mathbf{E}_{\text{pos}}\in\mathbb{R}^{(N+1)\times D}$ is added. Both the class token as well as the positional embeddings are learnable embeddings, which are trained jointly with the rest of the network. The resulting input sequence presented to the transformer network can be expressed as
$\mathbf{z}_{0}=[\mathbf{x}_{\text{class}};\thinspace\mathbf{x}_{p}^{1}\mathbf{E};\thinspace\mathbf{x}_{p}^{2}\mathbf{E};\thinspace\dots;\thinspace\mathbf{x}_{p}^{N}\mathbf{E}]+\mathbf{E}_{\text{pos}}.$
This sequence is presented to a standard transformer architecture of stacked attention modules. Each attention module consists of a multi-head self-attention (MSA) layer and a feedforward layer or multilayer perceptron (MLP) layer. MSA layers utilize self-attention (SA) (Vaswani et al., 2017), a powerful concept that allows transformers to relate and combine its feature embeddings with each other. Self-attention extracts features from the input sequence $\mathbf{z}$, which in turn preforms a transformation of the input vector sequence.
Specifically, self-attention extracts query, key and value sequences $\mathbf{q}$, $\mathbf{k}$ and $\mathbf{v}$ from the input sequence using a linear projection $\mathbf{U}_{qkv}\in\mathbb{R}^{D\times 3D_{h}}$: $[\mathbf{q},\mathbf{k},\mathbf{v}]=\mathbf{z}\mathbf{U}_{qkv}.$ The $\mathbf{q}$ and $\mathbf{k}$ sequences are then used to compute a Softmax-normalized transformation matrix $\mathbf{A}$ indicating how to incorporate information of the whole sequence (i.e. in our case all image patches) for every single vector of the sequence: $\mathbf{A}=\text{Softmax}(\frac{\mathbf{q}\mathbf{k}^{\intercal}}{\sqrt{D_{h}}}).$ Scaling the dot-product product by $\sqrt{D_{h}}$ here ensures a balanced distribution of the Softmax output. After obtaining $\mathbf{A}$, the output of SA is computed as $\text{SA}(\mathbf{z})=\mathbf{A}\mathbf{v}.$
A multi-head self-attention (MSA) layer (Vaswani et al., 2017) performs several attention operations in parallel, concatenates the result and projects it back to the internally used latent dimension of the transformer:
$\text{MSA}(\mathbf{z})=[\text{SA}_{1}(\mathbf{z});\thinspace\text{SA}_{2}(\mathbf{z});\thinspace\dots;\thinspace\text{SA}_{k}(\mathbf{z})]\mathbf{U}_{\text{msa}}$
In an attention module the multi-head self-attention layer is followed by a multi-layer-perceptron (MLP) layer transforming the recently combined embeddings to extract new feature representations. Before presenting $\mathbf{z}$ to each layer in the module, the embeddings are normalized using LayerNorm (Ba et al., 2016). To ensure consistent gradient flow during training, residual connections (He et al., 2016) are behind both the MSA and the MLP layers (Wang et al., 2019). Furthermore, as a regularization measure, Dropout (Srivastava et al., 2014) is applied after every MSA and MLP layer. In summary, given the sequence $\mathbf{z}_{t-1}$ from a previous attention module as input, we first compute the intermediate representation
$\mathbf{z}_{t}^{\prime}=\text{MSA}(\text{LN}(\mathbf{z}_{t-1}))+\mathbf{z}_{t-1},$ (1)
which is the presented to the MLP layer to compute the final output of the module
$\mathbf{z}_{t}=\text{MLP}(\text{LN}(\mathbf{z}_{t}^{\prime}))+\mathbf{z}_{t}^{\prime}.$ (2)
Finally, after N attention modules, the first vector of the sequence (corresponding to the class-token in the preprocessed input) is handed to a linear layer $\mathbf{W}_{\text{out}}\in\mathbb{R}^{D\times C}$ to predict the final class of the image: $\mathbf{y}=\text{argmax}(\mathbf{z}_{L}^{0}\mathbf{W}_{\text{out}}).$ $C$ denotes the number of classes.
@@ 3.2 Randomly permuting layer order during forward pass
During each forward pass, i.e. for each batch presented to the vision transformer, we randomly permute the execution order of layers during training. The intention here is to teach the layers to not only extract meaningful intermediate representations when receiving input from a particular layer, but to be able to process and encode information from and for all possible layers in the network. In terms of training, exchanging the order of layers does not require any changes in the basic error backpropagation algorithm. For the forward path, the order how weight matrices are multiplied and activation and attention functions applied changes for every batch and forward pass. This needs to be accounted in the backward pass by propagating the gradients in the precise reverse order that has been set in the forward pass, i.e.  multiplying the computed per-layer gradient matrices in the correct order. As we use Pytorch (Paszke et al., 2019) in all our experiments, this aspect is taken care of the framework’s autogradient feature. We refer to this model as LayerShuffle. To further illustrate the approach, a pseudocode listing is given in Algorithm 1.
@@
  Return $\mathbf{z}_{L}$ to be postprocessed by the transformer’s output layer.
@@ 4 Experiments
We conduct our experiments on the ILSVRC2012 dataset (Russakovsky et al., 2015), more commonly termed ImageNet2012. We use the original ViT-B/16 (Dosovitskiy et al., 2020) vision transformer, with the publicly available pre-trained weights provided on huggingface (vit, ). This model has been pre-trained on ImageNet21k (Ridnik et al., 2021) at an 224$\times$224 input image resolution and refined on ImageNet2012 at the same resolution. We refine the model again on ImageNet2012 at the same resolution, but using the training processes as described in Section 3. That is, we randomly permute layer execution order while refining the model. To establish a baseline, we refine the original ViT-B/16 for one more epoch without changing the layer order. We have limited ourselves to a single epoch here as we found any longer training to be unlikely to bring additional improvement in preliminary experiments. For each approach, including the baseline, we train $5$ networks and compare their average validation accuracy.
All models are refined using Adam (Kingma & Ba, 2014) ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-6}$), where an initial learning rate of $10^{-4}$ was empirically found to work best. In terms of batch size, we evaluate training batch sizes of 640 images, which is the maximum multiple of 8 that can fit in the video memory of our used GPU, as well as 128 images for models that benefit from a smaller batch size. We find that even smaller batch sizes do not yield any improvement in performance for our models. Inspecting training curves shows that the performance of our models plateaus at 20 epochs the latest, which we therefore set as the maximum number of training epochs. We use a form of early stopping by evaluating the model achieving lowest crossentropy loss on the validation set after 20 epochs. All models have been trained on a single NVIDIA H100 Tensor Core GPU with 80GB of memory, where the training of a single model for 20 epochs takes about 7 hours.
@@
We report the average accuracy for the baseline and LayerShuffle approaches in Table 1. Our baseline refined from a pre-trained ViT-B/16 model performs very much as expected: For a classic sequential execution order of the model layers, the average validation accuracy performs very close to the original model (Dosovitskiy et al., 2020) at $82.61\%$ at a low standard deviation of 0.08. Not surprisingly, for an arbitrary layer execution order, the average model accuracy declines catastrophically to $0.13\%$. Our original assertion that in general, neural networks are not robust to changing the execution order of their layers, is in line with these results.
Our LayerShuffle approach shows slightly lower performance than the baseline when executing layers in their original order (average accuracies of $75.22$, $75.28$, and $74.41$ for our LayerShuffle, LayerShuffle-position and LayerShuffle-predict respectively), but improves dramatically over the baseline model in an arbitrary execution order setting. Surprisingly, the simplest LayerShuffle approach receives an accuracy of $62.77\%$, only slightly lower than the accuracy 63.61% of the LayerShuffle-position, which receives information about the layer position. While all three approaches seem to be overall on par, we observe the position prediction approach, LayerShuffle-predict, to be the least well-performing of the proposed approaches with an average accuracy of $61.18\%$. A possible explanation might be that due to optimization of multiple objectives (fitting both the output labels as well as predicting the current position of the layer) this approach requires more careful hyperparameter tuning. This hypothesis is also supported by the somewhat higher standard deviation of $1.06$ percentage points in accuracy, compared to the $0.41$ and $0.23$ percentage points of LayerShuffle and LayerShuffle-position respectively. A further interesting observation is to be made when comparing the performance for sequential and arbitrary execution order for each approach respectively. For all approaches, using the original layer order for sequential execution still performs better than an arbitrary order. This is most likely a consequence of fine-tuning from a sequentially trained model.
@@
For the layer position prediction approach, we measure the average accuracy of layer position predictions over all five trained LayerShuffle-predict models, and find that the layer position is predicted correctly in $99.99\%$ of all cases. These results demonstrate that each layer has enough information coming from its inputs alone to predict where it is in the network, providing the basis to adapt to its current position. We investigate this further when analyzing intermediate network representations in Section 4.3.
In conclusion, refining a pre-trained model while randomly permuting the execution order of the network layers can make a model more robust towards such arbitrary execution orders at test time. On the other hand, Dropout and LayerNorm by themselves do not have the same effect and fail to produce networks robust against layer shuffling.
@@
Table 1: Approach accuracy for sequential and arbitrary execution order of layers on the ImageNet2012 validation set. The baseline model (ViT-B/16) performs best when executed sequentially but fails catastrophically when the layers are executed in an arbitrary order. Even the simplest LayerShuffle variant (LayerShuffle), in which the model does not have any information about its current position, reaches an accuracy of 62.77. All of our proposed training approaches permit the models to be executed with arbitrary layer execution order at test time, while still delivering good performance for the original model execution order.
layer order ViT-B/16 LayerShuffle LayerShuffle-position LayerShuffle-predict
sequential $\mathbf{82.61\pm 0.08\%}$ $75.22\pm 0.28\%$ $\mathit{75.28\pm 0.18\%}$ $74.41\pm 0.20\%$
arbitrary $0.13\pm 0.03\%$ $\mathit{62.77\pm 0.41\%}$ $\mathbf{63.61\pm 0.23\%}$ $61.18\pm 1.06\%$

### v2

@@ Abstract
Due to their architecture and how they are trained, artificial neural networks are typically not robust toward pruning or shuffling layers at test time. However, such properties would be desirable for different applications, such as distributed neural network architectures where the order of execution cannot be guaranteed or parts of the network can fail during inference. In this work, we address these issues through a number of training approaches for vision transformers whose most important component is randomizing the execution order of attention modules at training time. With our proposed approaches, vision transformers are capable to adapt to arbitrary layer execution orders at test time assuming one tolerates a reduction (about 20%) in accuracy at the same model size. We analyse the feature representations of our trained models as well as how each layer contributes to the models prediction based on its position during inference. Our analysis shows that layers learn to contribute differently based on their position in the network. Finally, we layer-prune our models at test time and find that their performance declines gracefully. Code available at https://github.com/matfrei/layershuffle.
@@ 1 Introduction
While demonstrating impressive performance in many domains (Krizhevsky et al., 2012; Vaswani et al., 2017; Radford et al., 2021; Rombach et al., 2022), deep learning systems demand both extensive computational resources and tight integration of their parts. For applications at scale, they therefore increasingly require the construction of large data centers with thousands of dedicated hardware accelerators. A paradigm shift from central to decentral model inference, where loosely coupled neural networks are distributed over a number of edge devices that share the computational load of the model (Gacoin et al., 2019) therefore seems ultimately desirable. Unfortunately, current deep learning models lack the robustness necessary for such a paradigm shift.
In general, artificial neural networks (ANNs) are not robust toward pruning or replacing network layers during deployment. Similarly, changing the order of execution in-between layers without further training usually results in catastrophic losses in accuracy. Nevertheless, these properties would be desirable e.g. in distributed setups as described above, where a model is executed on a number of shared nodes in a network. This way, overloaded or malfunctioning nodes could simply be skipped in favor of other available nodes.
@@
The more recently discovered transformer architecture (Vaswani et al., 2017) has been shown to be more flexible. Transformers, when trained accordingly, can be layer-pruned at test-time (Fan et al., 2019), and recent work merges similar transformer-based language models (Akiba et al., 2024), all with only moderate reduction or even an improvement in performance. We hypothesize that the reason for the high adaptability of transformers can be found in self-attention modules being able to adapt their output based on the received input. Thus it should be possible to train a transformer network to not only adapt to the variation of its input features based on the overall network input but also the variations caused by receiving input from different layers during test time.
We propose and evaluate three training approaches for vision transformers to address the robustness issues laid out above. The most important component common to all approaches is randomizing the execution order of the vision transformer’s stacked self-attention-and-feed-forward modules at training time (Figure 1(a)). More precisely, the main contributions in this paper are:
  • With LayerShuffle, the layers of a vision transformer (Dosovitskiy et al., 2020) are capable of adapting to an arbitrary execution order at test time, assuming one tolerates a moderate reduction in performance. Providing each layer additionally with its current position in the network improves performance only slightly compared to a model without it, suggesting that each attention layer is already capable of determining its role based on the incoming data alone.
  • An analysis reveals that layers of models trained with LayerShuffle adjust their output depending on which position they hold in the network.
  • Trained models can be layer-pruned at test time similar to the models trained with the techniques proposed in Fan et al. (2019), where their performance declines gracefully, i.e. models with reduced amounts of layers still remain functional.
@@ 2 Related work
Zhu et al. (2020) find that for particular subsets of inputs, transformers perform better when changing the execution order of layers to an input-specific sequence. They optimize the execution order per sample in order to maximize the performance of the model for natural language processing tasks.
While the goal in their work is to find a layer sequence of a pre-trained model that is optimal for a given input, our approach aims to make the model robust to any sequence of execution, where layers might even be missing.
In parallel to our work on vision transformers, two groups have conducted similar experiments with the aim to understand how language models (LLMs) process data. Lad et al. (2024) found that LLMs are very robust to changing the positions of adjacent layers or ablating single layers from the model. Sun et al. (2024) perform similar experiments, and find that transformers improve iteratively upon their predictive output by subsequently refining the internal representation of the presented input. The main difference to our work is that the authors of these works do not perform any refinement on the models and switch and ablate layers locally with the aim of better understanding the inner workings of LLMs. Here we focus on methods and training approaches to increase this innate robustness of the transformer architecture to a point where models at test time function regardless of their layer execution order, and respond gracefully to the ablation of several layers in any position of the network.
@@
Work on introducing permutation invariance into neural networks has been conducted by Lee et al. (2019), Tang & Ha (2021) as well as Pedersen & Risi (2022). The corresponding former two approaches exploit the permutation equivariance of attention, i.e. the fact that the order in which a sequence of vectors gets presented to the attention module does not change its result, but merely shuffles the sequence of output vectors. This equivariance is achieved by using a fixed-seed query vector in order to obtain an permutation invariant latent code. This latent code stays the same no matter in which order input tokens/patches are presented to the module. The main contrast to our work here is that we exploit permutation invariance in the order of layer executions rather than the order of tokens and patch embeddings and can therefore not make use of permutation equivariance of the attention operation, as it does not apply to switching inputs and outputs.
@@ 3 Methods
We investigate three approaches for arbitrary layer execution order in vision transformers (ViT; Dosovitskiy et al., 2020): First, we simply permute the order of layers randomly during training, such that every training batch is presented to the network’s layers in a different random order (Section 3.1). Second, while randomly permuting the layer order as in the previous approach, we use an layer-depth encoding inspired by learned word embedding approaches (Section 3.2) to test if this additional information would further improve performance. Third, while randomly permuting layer order as in the previous approaches, we try to predict from the output of every layer at which position the layer is currently located in the network using a small layer position prediction network for every layer (Section 3.3). A detailed overview on ViTs can be found in the appendix.
@@ 3.1 Randomly permuting layer order during forward pass
During each forward pass, i.e. for each batch presented to the ViT, we randomly permute the execution order of layers during training. The intention here is to teach the layers to not only extract meaningful intermediate representations when receiving input from a particular layer, but to be able to process and encode information from and for all possible layers in the network. In terms of training, exchanging the order of layers does not require any changes in the basic error backpropagation algorithm. For the forward path, the order how weight matrices are multiplied and activation and attention functions applied changes for every batch and forward pass. This needs to be accounted in the backward pass by propagating the gradients in the precise reverse order that has been set in the forward pass, i.e.  multiplying the computed per-layer gradient matrices in the correct order. As we use Pytorch (Paszke et al., 2019) in all our experiments, this aspect is taken care of the framework’s autogradient feature. We refer to this model as LayerShuffle. To further illustrate the approach, a pseudo-code listing is given in Algorithm 1.
@@
  Return $\mathbf{z}_{L}$ to be post-processed by the transformer’s output layer.
@@ 4 Experiments
We conduct our experiments on the ILSVRC2012 dataset (Russakovsky et al., 2015), more commonly termed ImageNet2012, as well as the CIFAR-100 dataset Krizhevsky et al. . We use the original ViT-B/16 (Dosovitskiy et al., 2020) vision transformer, as well the DeiT-B distilled data-efficient image transformer Touvron et al. (2021). Pre-trained weights for ImageNet2012 are publicly available for both models (Dosovitskiy et al., 2020; Wu et al., 2020; Touvron et al., 2021).
The ViT-B/16 has been pre-trained on ImageNet21k (Deng et al., 2009; Ridnik et al., 2021) at an 224$\times$224 input image resolution and refined on ImageNet2012 at the same resolution. DeiT-B has the same architecture as ViT-B/16, but uses an additional destillation token during training, which is used to distill the inductive bias of a large convolutional network into the transformer in order to require less training data. It is pre-trained exclusively on ImageNet2012.
Both models are again refined on both ImageNet2012 and CIFAR-100 at the same resolution, but using the training processes as described in Section 3. That is, layer execution order is randomly permuted while refining the model. To establish a baseline, on ImageNet2012, we refine the original models for one more epoch on without changing the layer order. Any longer training was found unlikely to bring additional improvement in preliminary experiments since our networks are already pretrained on ImageNet. For CIFAR-100, we refine our baseline for 20 epochs. For each approach, including the baselines, we train $5$ networks and compare their average validation accuracy.
All models are refined using Adam (Kingma & Ba, 2014) ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-6}$), where an initial learning rate of $10^{-4}$ was empirically found to work best. In terms of batch size, we evaluate training batch sizes of 640 images, which is the maximum multiple of 8 that can fit in the video memory of our used GPU, as well as 128 images for models that benefit from a smaller batch size. Even smaller batch sizes do not yield any improvement in performance for our models. Inspecting training curves shows that for ImageNet2012 the performance of models plateaus at 20 epochs the latest, which is therefore set as the maximum number of training epochs. For CIFAR-100 we use 100 epochs since the models were pretrained on a different dataset, i.e. ImageNet. We use a form of early stopping by evaluating the model achieving the lowest crossentropy loss on the validation set after the maximum amount of training epochs. All models have been trained on a single NVIDIA H100 Tensor Core GPU with 80GB of memory. Training a single model on ImageNet2012 for 20 epochs takes about 7 hours whereas CIFAR-100 training times are significantly shorter due to the smaller train set.
@@
The average accuracy for all approach on all vision transformer architectures and datasets is shown in Table 1. We make the following observations across all models and datasets:
@@
On both the CIFAR-100 and the ImageNet2012 dataset, our baselines refined from pre-trained ViT-B/16 and DeiT-B models perform very much as expected. For a classic sequential execution order of the model layers, on ImageNet2012 the trained models achieve an average validation accuracy very close to the performance of the respective original pretrained models (Dosovitskiy et al., 2020; Touvron et al., 2021). Our refined baseline ViT-B/16 obtains an average accuracy of $82.61\%$ with a standard deviation of $0.08$. The DeiT-B model attains a slightly lower accuracy of $81.16\%$ with a standard deviation of $0.06$. Baseline results CIFAR-100 look similar with ViT-B/16 and DeiT-B achieving $89.62\%$ (standard deviation: $0.26$) and $87.31\%$ (standard deviation: $0.25$) respectively. Not surprisingly, for an arbitrary layer execution order, the average model accuracy declines catastrophically to below $1\%$ for all trained models on both datasets. Our original assertion that in general, ANNs are not robust to changing the execution order of their layers, is in line with these results.
Our LayerShuffle approaches show lower performance than the baselines when executing layers in their original order. On ImageNet2012, our trained ViT-B/16 models obtain average accuracies of $75.22$, $75.28$, and $74.41$ respectively for our LayerShuffle, LayerShuffle-position and LayerShuffle-predict approaches. Average accuracies for DeiT-B models are in a similar range with $76.57$, $76.18$, and $75.08$ for our respective LayerShuffle, LayerShuffle-position and LayerShuffle-predict approaches. On CIFAR-100 on the other hand the gap between baseline and LayerShuffle models is somewhat larger for the trained ViT-B/16 models. These models merely achieve $64.43$, $61.98$ and $60.88$ with slightly higher standard deviations for LayerShuffle, LayerShuffle-position and LayerShuffle-predict approaches. For DeiT-B models on the other hand, these approaches perform similarly well to the models trained on ImageNet with scores of $75.04$, $72.13$ and $66.94$ for the above mentioned techniques. A possible explanation for these discrepancies could be found in ViT-B/16 models requiring more and more diverse training data compared to DeiT-B models, which has been pre-trained with the aim to reduce the amount of required training data.
Despite being outperformed by the baseline in a sequential execution order setting, all models improve dramatically over their corresponding baseline models in an arbitrary execution order setting. Taking a closer look at LayerShuffle model performance in that setting, we find that the simplest approach performs very well across both architectures and datasets. For both ImageNet2012 as well as CIFAR-100 validation sets DeiT-B trained on LayerShuffle yields the best performance with average accuracies of $66.62$ and $64.99$ respectively, narrowly outperforming LayerShuffle-position, which receives information about the layer position. LayerShuffle-position achieves scores of $66.62$ and $64.99$ on these datasets. For ViT-B/16 models on the other hand, LayerShuffle-position outperforms LayerShuffle. The former achieves scores of $63.61$ and $55.47$ of ImageNet2012 and CIFAR-100, with the latter performing only slightly worse with $62.77$ and $54.34$. The most likely explanation for models of the ViT-B/16 architecture achieving significantly lower accuracies on CIFAR-100 than DeiT-B models can again be found in the former requiring less training data than the latter.
We find that the position prediction approach, LayerShuffle-predict is outperformed by both our remaining approaches on all datasets and architectures. On ImageNet2012, refined ViT-B/16 models achieve average accuracies of $61.18$ whereas DeiT-B models attain $64.51$. On CIFAR-100 the former score $53.53$, the latter $58.77$. A possible explanation might be that due to optimization of multiple objectives (fitting both the output labels as well as predicting the current position of the layer) this approach requires more careful hyperparameter tuning.
A further interesting observation is to be made when comparing the performance for sequential and arbitrary execution order for each approach respectively. For all approaches, using the original layer order for sequential execution still performs better than an arbitrary order. This is most likely a consequence of fine-tuning from a sequentially trained model.
@@
For the layer position prediction approach, we measure the average accuracy of layer position predictions over all five trained LayerShuffle-predict models, and find that the layer position is predicted correctly in $99.99\%$ of all cases. These results demonstrate that each layer has enough information coming from its inputs alone to predict where it is in the network, providing the basis to adapt to its current position. We investigate this further when analyzing intermediate network representations in Section 4.3. In conclusion, refining a pre-trained model while randomly permuting the execution order of the network layers can make a model more robust towards such arbitrary execution orders at test time. On the other hand, Dropout and LayerNorm by themselves do not have the same effect and fail to produce networks robust against layer shuffling.
@@
Table 1: Approach accuracy for sequential and arbitrary execution order of layers on the ImageNet2012 (IN2012) and CIFAR-100 (CIFAR) validation sets. The baseline models perform best when executed sequentially but fail catastrophically when the layers are executed in an arbitrary order. Even the simplest LayerShuffle variant, in which the model does not have any information about its current position, reaches accuracies above $60$ percent. All of our proposed training approaches permit the models to be executed with arbitrary layer execution order at test time, while still delivering good performance for the original model execution order.
model layer order baseline LS LS-pos LS-pred
IN2012 ViT-B/16 sequential $\mathbf{82.61}$±$\pm 0.08$ $75.22$±$\pm 0.28$ $\mathit{75.28}$±$\pm 0.18$ $74.41$±$\pm 0.20$
IN2012 ViT-B/16 arbitrary $0.13$±$\pm 0.03$ $\mathit{62.77}$±$\pm 0.41$ $\mathbf{63.61}$±$\pm 0.23$ $61.18$±$\pm 1.06$
IN2012 DeiT-B dist. sequential $\mathbf{81.16}$±$\pm 0.06$ $\mathit{76.57}$±$\pm 0.12$ $76.18$±$\pm 0.46$ $75.08$±$\pm 0.27$
IN2012 DeiT-B dist. arbitrary $0.12$±$\pm 0.05$ $\mathbf{66.62}$±$\pm 0.4$ $\mathit{65.89}$±$\pm 1.28$ $64.51$±$\pm 0.85$
CIFAR ViT-B/16 sequential $\mathbf{89.62}$±$\pm 0.26$ $\mathit{64.43}$±$\pm 0.95$ $61.98$±$\pm 0.49$ $60.88$±$\pm 0.76$
CIFAR ViT-B/16 arbitrary $0.64$±$\pm 0.13$ $\mathit{54.34}$±$\pm 2.64$ $\mathbf{55.47}$±$\pm 1.41$ $53.53$±$\pm 1.45$
CIFAR DeiT-B dist. sequential $\mathbf{87.31}$±$\pm 0.25$ $\mathit{75.04}$±$\pm 0.83$ $72.13$±$\pm 1.5$ $66.94$±$\pm 0.6$
CIFAR DeiT-B dist. arbitrary $0.53$±$\pm 0.16$ $\mathbf{64.99}$±$\pm 0.76$ $\mathit{64.31}$±$\pm 1.70$ $58.77$±$\pm 1.18$